OpenStudy (anonymous):

Diana has a box containing 6 blue, 8 red, 5 purple, 9 green, and 2 clear marbles that are all the same size and shape. What is the probability of randomly choosing a clear marble on the first pick, replacing it, and then randomly choosing a purple marble on the second pick? A- 1/6 B- 1/15 C- 1/87 D- 1/90

OpenStudy (sandra):

since these are independent events (and you place the first marble back), the answer is : 1. P(drawing a clear) * P(drawing a purple) 2. there are 30 marbles, and 2 clear marbles, the probability of drawing a clear is 2/30 3. once you put one back, there are still 30 marbles, and 5 purple marbles, and the probability of drawing a purple is 5/30 4. so the answer is 2/30 * 5/30 = 10/900 = 1/90